# Identifying Trustworthiness Challenges in Deep Learning Models for Continental-Scale Water Quality Prediction

### v1

@@ Abstract
Water quality is foundational to environmental sustainability, ecosystem resilience, and public health. Deep learning models, particularly Long Short-Term Memory (LSTM) networks, offer transformative potential for large-scale water quality prediction and scientific insights generation. However, their widespread adoption in high-stakes decision-making, such as pollution mitigation and equitable resource allocation, is prevented by unresolved trustworthiness challenges including fairness, uncertainty, interpretability, robustness, generalizability, and reproducibility. In this work, we present the first comprehensive evaluation of trustworthiness in a continental-scale multi-task LSTM model predicting 20 water quality variables (encompassing physical/chemical processes, geochemical weathering, and nutrient cycling) across 482 U.S. basins. Our investigation uncovers systematic patterns of model performance disparities linked to basin characteristics, the inherent complexity of biogeochemical processes, and variable predictability, emphasizing critical performance fairness concerns. We further propose methodological frameworks for quantitatively evaluating critical aspects of trustworthiness, including uncertainty, interpretability, and robustness, identifying key limitations that could challenge reliable real-world deployment. This work serves as a timely call to action for advancing trustworthy data-driven methods for water resources management and provides a pathway to offering critical insights for researchers, decision-makers, and practitioners seeking to leverage artificial intelligence (AI) responsibly in environmental management.
1 Introduction
@@
Water quality is essential for both environmental sustainability and public health, and clean water supports aquatic biodiversity, ecosystem services, and safe drinking water access [1, 2]. However, freshwater contaminants increasingly threaten ecological integrity and human well-being by causing habitat degradation, species loss, and waterborne disease [3]. In response, substantial efforts have been made, notably with the U.S. investing more than $1.9 trillion since 1960 to reduce pollution in rivers, lakes, and other surface waters, a commitment exceeding most other U.S. environmental initiatives [4]. Assessment of water quality is typically based on sample collection that is expensive and labor-intensive, while also producing highly heterogeneous datasets with limited scalability [5, 6].
@@
Artificial Intelligence (AI) has provided a new set of tools with the potential to make progress on these challenges. In particular, deep neural networks such as long short-term memory (LSTM) models, renowned for their ability to capture temporal dependencies in time series data, have become the most popular machine learning models in hydrological modeling [7, 8, 9]. However, their application to predict a wide range of water quality variables over space and time, especially across large scales, remains nascent and contentious. Practitioners often distrust “black-box” machine learning frameworks, despite improved predictive accuracy relative to simple statistical models, due to concerns about explainability and interpretability [10]. This skepticism arises partly because of the high stakes involved: water quality management decisions carry profound societal implications, from infrastructure design to pollution mitigation. To be adopted and used, AI-based models need to be not only accurate but also trustworthy.
@@
Trustworthiness in AI broadly encompasses multiple critical aspects, including but not limited to their fairness in both modeling processes and outcomes, uncertainty quantification, explainability of model decisions, robustness against noise and adversarial disruption, generalizability to unseen conditions, and reproducibility of research [11, 12, 13, 14, 15]. While building trustworthy data-driven methods has long been a vision and key development goal in many fields of science and engineering [16], such as healthcare, autonomous driving, sentiment analysis, and climate science [17, 18, 19, 20], their adoption in water quality research lags considerably behind. Current deep learning research in water quality has largely focused on improving performance metrics (e.g., predictive accuracy), often overlooking these essential trustworthiness considerations. For deep learning models to become more widely adopted in guiding watershed management and decision-making, especially when combined with climate and socioeconomic uncertainties, multiple aspects of trustworthiness need to be more carefully evaluated and justified [21].
@@
In this study, we present the first comprehensive evaluation of trustworthiness in LSTM-based prediction of water quality variables. This investigation extends beyond conventional predictive accuracy metrics to address persistent concerns: (1) How does predictive accuracy vary across water quality variables and basin types? (2) How does uncertainty in model predictions vary across variables, and are certain variables disproportionately impacted? (3) Do feature importance analysis methods yield consistent results in identifying key drivers? (4) How robust is the model when exposed to outliers, random noise, and adversarial perturbations? In addition, by advocating for generalizability and reproducibility, we aim to emphasize the significance of advancing trustworthy data-driven methods for water quality management and provide a pathway to offering critical insights for researchers, policymakers, and practitioners seeking to leverage AI responsibly in environmental management. Our analysis makes use of datasets covering 37 years of hydroclimate forcings, rainfall chemistry, vegetation indices, and static basin attributes across 482 U.S. basins (see Methods). We train a multi-task LSTM model to predict 20 water quality variables associated with physical/chemical, geochemical weathering, and nutrient cycling processes. We employ the Monte Carlo dropout approach [22] to assess prediction uncertainty, compare feature importance derived from ablation studies and the Traverse method, and assess the robustness of LSTM-based models under various data corruption scenarios.
2 Results and Discussion
2.1 Challenges in predicting management-critical water quality variables
@@
2.2 Fairness concerns due to land use heterogeneity, data availability, and variable simplicity
@@
2.3 Higher prediction uncertainty in management-critical water quality variables
@@
2.4 Inconsistent feature importances challenge model interpretability
@@
2.5 Limited robustness to outliers, random noise, and adversarial disruptions
Robustness against random noise. Similar to the effects observed in the outlier injection experiments, our results also show that LSTM models are sensitive to random noise, with varying degrees of impact across different water quality variables (Fig. 4c). Likewise, the performance degradation is significantly higher when noise is applied to target water quality variables compared to input features (), except for TSS, OrgN, , and TP. Strong correlations between the baseline median KGE and the percentage change of KGE ( for input features and for water quality variables, Fig. 4d) suggest that enhancing baseline predictive accuracy could improve robustness to random noise in both input features and target variables.
While this study focuses on LSTM architectures to benchmark robustness challenges in water quality research, it also highlights the need for the broader community to investigate whether these vulnerabilities are unique to LSTMs or inherent to deep learning approaches more broadly. Future research should also compare LSTMs with alternative deep learning architectures, such as Transformer-based models [37], to disentangle limitations associated with sequential modeling from fundamental deep learning vulnerabilities (e.g., sensitivity to input perturbations, overreliance on training distributions). For instance, Transformer models, which replace recurrent connections with self-attention mechanisms, may exhibit distinct robustness due to their ability to capture any long-term pair-wised dependencies and parallelize training. Such comparisons may be able to provide insights into whether architectural innovations can help mitigate vulnerabilities to outliers, random noise, and adversarial attacks.
2.6 Generalizability
@@
Another persistent challenge for deep learning models in water quality prediction lies in their limited capacity to generalize beyond training conditions. Unlike process-based models that embed mechanistic relationships (e.g., reaction kinetics and advection-diffusion), LSTMs rely on statistical patterns in training data, which are often sparsely measured for critical water quality variables or inadequately capture extreme hydrological events such as floods and droughts [38]. This limitation is exacerbated by climate change, which amplifies the frequency and magnitude of extremes that lie outside historical training distributions [39].
@@
Knowledge-Informed Deep Learning (KIDL) has emerged as a promising strategy to bridge this gap. By incorporating domain knowledge, such as mass conservation and transport dynamics into the model’s architecture and learning process, KIDL enforces physically plausible behavior even under unseen conditions. For example, Agrawal et al. [40] integrated a physical “flow-gate” mechanism into an LSTM model to explicitly model hysteresis between discharge and solute dynamics, which improved predictions of nine stream solutes (RMSE reduced by 1-32% compared with standard LSTM). Hybrid modeling approaches have also shown success by combining physics-based simulations with physics-informed loss functions to ensure predictions align with foundational physical principles, as demonstrated in lake temperature modeling [41]. While these studies illustrate the potential of KIDL, its application in water quality research remains limited, compared to its use in hydrology and other scientific fields [42, 43, 44].
In addition, large-scale pretrained models offer complementary pathways to enhance generalizability. These models have the ability to transfer learned knowledge to unfamiliar tasks via few-shot (minimal samples) or zero-shot (no samples) learning [45]. For example, a model pretrained on extensive, diverse datasets, even those unrelated specifically to water quality, may effectively generalize to predict water quality variables in ungauged basins using limited local measurements. By decreasing dependence on sparse, domain-specific data, pretrained models have the potential to substantially improve resilience to distribution shifts driven by climate change.
2.7 Reproducibility
@@
3 Methods
3.1 Water quality data and basin selection
@@ 3.2 Model input
3.3 Multi-task LSTM model training and evaluation
The Long Short-Term Memory (LSTM) model is a prominent member of Recurrent Neural Network (RNN) models designed to leverage sequential information for time series prediction [54]. Unlike standard RNNs, which suffer from the vanishing gradient problem when capturing long-term dependencies [55]. LSTM incorporates a memory mechanism to address this limitation. This mechanism, involving “memory states” and “gates”, allows the model to regulate what information to retain or discard over time, enabling more effective learning of temporal patterns. In this work, we trained a multi-task LSTM model to predict 20 water quality variables simultaneously. Unlike single-task models that require training separate models for each target variable [29, 52, 56], the multi-task approach leverages shared learning of hidden relationships among variables, improving predictive performance, and reducing computational costs [23], which is critical for modeling complex biogeochemical processes.
@@ 3.3 Multi-task LSTM model training and evaluation
3.4 Trustworthiness evaluation framework

### v2

@@ Abstract
Water quality is foundational to environmental sustainability, ecosystem resilience, and public health. Deep learning models, particularly Long Short-Term Memory (LSTM) networks, offer transformative potential for large-scale water quality prediction and scientific insights generation. However, their widespread adoption in high-stakes decision-making, such as pollution mitigation and equitable resource allocation, is prevented by unresolved trustworthiness challenges including fairness, robustness, uncertainty, interpretability, generalizability, and reproducibility. In this work, we present the first comprehensive, quantitative evaluation of trustworthiness in a continental-scale multi-task LSTM model predicting 20 water quality variables (encompassing physical/chemical processes, geochemical weathering, and nutrient cycling) across 482 U.S. basins. Our investigation uncovers systematic patterns of model performance disparities linked to basin characteristics, the inherent complexity of biogeochemical processes, and variable predictability, emphasizing critical performance fairness concerns. We further propose methodological frameworks for quantitatively evaluating critical aspects of trustworthiness, including robustness, uncertainty, and interpretability, identifying key limitations that could challenge reliable real-world deployment. This work serves as a timely call to action for advancing trustworthy data-driven methods for water resources management and provides a pathway to offering critical insights for researchers, decision-makers, and practitioners seeking to leverage artificial intelligence (AI) responsibly in environmental management.
Introduction
@@
Water quality is essential for both environmental sustainability and public health, and clean water supports aquatic biodiversity, ecosystem services, and safe drinking water access [1, 2]. However, freshwater contaminants increasingly threaten ecological integrity and human well-being by causing habitat degradation, species loss, and waterborne diseases [3]. In response, substantial efforts have been made, notably with the U.S. investing more than $1.9 trillion since 1960 to reduce pollution in rivers, lakes, and other surface waters, a commitment exceeding most other U.S. environmental initiatives [4]. Assessment of water quality is typically based on sample collection that is expensive and labor-intensive, while also producing highly heterogeneous datasets with limited scalability [5, 6].
@@
Artificial Intelligence (AI) has provided a new set of tools with the potential to make progress on these challenges. In particular, deep neural networks such as long short-term memory (LSTM) models, renowned for their ability to capture long temporal dependencies in time series data, have become the most widely applied machine learning models in hydrological modeling [7, 8, 9]. However, their application to predict a wide range of water quality variables over space and time, especially across large scales, remains nascent and contentious. In addition, practitioners often distrust “black-box” machine learning frameworks, despite improved predictive accuracy relative to simple statistical models, due to concerns about explainability and interpretability [10]. This skepticism arises partly because of the high stakes involved: water quality management decisions carry profound societal implications, from infrastructure design to pollution mitigation. To be adopted and used, AI-based models need to be not only accurate but also trustworthy.
@@
Trustworthiness in AI broadly encompasses multiple critical aspects, including but not limited to their fairness in both modeling processes and outcomes, robustness against noise and adversarial disruption, uncertainty quantification, explainability of model decisions, generalizability to unseen conditions, and reproducibility of research [11, 12, 13, 14, 15]. While building trustworthy data-driven methods has long been a vision and key development goal in many fields of science and engineering [16], such as healthcare, autonomous driving, sentiment analysis, and climate science [17, 18, 19, 20], their adoption in water quality research lags considerably behind. Current deep learning research in water quality has largely focused on improving performance metrics (e.g., predictive accuracy), often overlooking these essential trustworthiness considerations. For deep learning models to become more widely adopted in guiding watershed management and decision-making, especially when combined with climate and socioeconomic uncertainties, multiple aspects of trustworthiness need to be more carefully evaluated and justified [21].
@@
In this study, we present the first comprehensive, quantitative evaluation of trustworthiness in LSTM-based prediction of water quality variables. This investigation extends beyond conventional predictive accuracy metrics to address persistent concerns: (1) How does model performance vary across water quality variables and basin types? (2) How robust is the model when exposed to outliers, random noise, and adversarial perturbations? (3) How does uncertainty in model predictions vary across variables, and are certain variables disproportionately impacted? (4) Do feature importance analysis methods yield consistent results in identifying key drivers? (5) Can the model generalize effectively to unmonitored basins? To answer these questions, we first train a multi-task LSTM model using 37 years of water quality observations, hydroclimate forcings, and static basin attributes across 482 U.S. basins (see Methods) to predict 20 water quality variables representing physical/chemical, geochemical weathering, and nutrient cycling processes. We then evaluate the model’s robustness under various data corruption scenarios, quantify prediction uncertainty using the Monte Carlo dropout approach [22], and compare feature importance derived from ablation studies and the Traverse method. Our main contribution is defining and demonstrating a reproducible trust benchmarking protocol, which is urgently needed yet missing in this domain. The proposed evaluation framework is model-agnostic and can be readily extended to other architectures or application domains.
Results and Discussion
Challenges in predicting management-critical water quality variables
@@
Fairness concerns due to land use heterogeneity, data availability, and variable simplicity
@@
Limited robustness to outliers, random noise, and adversarial disruptions
Robustness against random noise. Similar to the effects observed in the outlier injection experiments, our results also show that LSTM models are sensitive to random noise, with varying degrees of impact across different water quality variables (Fig. 2c). Likewise, the performance degradation is significantly higher when noise is applied to target water quality variables compared to input features (), except for TSS, OrgN, , and TP. Strong correlations between the baseline median KGE and the percentage change of KGE ( for input features and for water quality variables, Fig. 2d) suggest that enhancing baseline predictive accuracy could improve robustness to random noise in input features and target variables.
While this study focuses on LSTM architectures to benchmark robustness challenges in water quality research, it also highlights the need for the broader community to investigate whether these vulnerabilities are unique to LSTMs or inherent to deep learning approaches more broadly. Future research should also compare LSTMs with alternative deep learning architectures, such as Transformer-based models [27], to disentangle limitations associated with sequential modeling from fundamental deep learning vulnerabilities (e.g., sensitivity to input perturbations, overreliance on training distributions). For instance, Transformer models, which replace recurrent connections with self-attention mechanisms, may exhibit distinct robustness due to their ability to capture any long-term pair-wise dependencies and parallelize training. Such comparisons may be able to provide insights into whether architectural innovations can help mitigate vulnerabilities to outliers, random noise, and adversarial attacks.
Higher prediction uncertainty in management-critical water quality variables
@@
Inconsistent feature importances challenge model interpretability
@@
Our study intentionally focused on two widely used and computationally feasible methods, Traverse and Ablation, that are effective for evaluating model trustworthiness. The goal is not to provide an exhaustive comparison of all interpretability techniques. While other methods such as SHAP [37] and Integrated Gradients [38] are valuable methods and are also used for feature importance analysis, their omission does not diminish the validity or relevance of the trust benchmarking framework we propose.
Generalizability
Our spatial training-testing split results (Fig. S6) indicate that the LSTM model exhibits poor generalizability to “ungauged” basins for most water quality variables, except for temperature and dissolved oxygen, which achieve relatively high KGE values of 0.88 and 0.71, respectively. These findings highlight the challenge of learning spatially transferable patterns, particularly for those geochemical and nutrient-related variables.
@@
The sparsity of water quality observations, with many basins having limited or no historical data, further compounds this challenge. In practice, a trustworthy model must generalize reasonably well to data-scarce basins, enabling reliable predictions where in situ monitoring is unavailable. However, a fundamental limitation of deep learning models lies in their limited capacity to generalize beyond training conditions. Unlike process-based models that embed mechanistic relationships (e.g., reaction kinetics and advection-diffusion), LSTMs rely on statistical patterns in training data, which are often sparsely measured for critical water quality variables or inadequately capture extreme hydrological events such as floods and droughts [39]. As climate change continuously increases the frequency and severity of such extremes [40], the inability of purely data-driven models to extrapolate beyond their training distribution poses a critical challenge for trustworthy water quality prediction.
@@
Knowledge-Informed Deep Learning (KIDL) has emerged as a promising strategy to address these challenges. By incorporating domain knowledge, such as mass conservation and transport dynamics into the model’s architecture and learning process, KIDL enforces physically plausible behavior even under unseen conditions. For example, Agrawal et al. [41] integrated a physical “flow-gate” mechanism into an LSTM model to explicitly model hysteresis between discharge and solute dynamics, which improved predictions of nine stream solutes (RMSE reduced by 1-32% compared with standard LSTM). Hybrid modeling approaches have also shown success by combining physics-based simulations with physics-informed loss functions to ensure predictions align with foundational physical principles, as demonstrated in lake temperature modeling [42]. While these studies illustrate the potential of KIDL, its application in water quality research remains limited, compared to its use in hydrology and other scientific fields [43, 44, 45]. In addition, large-scale pretrained foundation models offer complementary pathways to enhance generalizability. These models have the ability to transfer learned knowledge to unseen tasks via few-shot (minimal samples) or zero-shot (no samples) learning [46]. For example, a model pretrained on extensive, diverse datasets, even those unrelated specifically to water quality, may effectively generalize to predict water quality variables in ungauged basins using limited local measurements. By decreasing dependence on sparse, domain-specific data, pretrained models have the potential to substantially improve resilience to distribution shifts driven by climate change.
Reproducibility
@@
Methods
Water quality data and basin selection
@@ Methods
Multi-task LSTM model training and evaluation
The Long Short-Term Memory (LSTM) model is a prominent member of Recurrent Neural Network (RNN) models designed to leverage sequential information for time series prediction [55]. Unlike standard RNNs, which suffer from the vanishing gradient problem when capturing long-term dependencies [56]. LSTM incorporates a memory mechanism to address this limitation. This mechanism, involving “memory states” and “gates”, allows the model to regulate what information to retain or discard over time, enabling more effective learning of temporal patterns. In this work, we trained a multi-task LSTM model to predict 20 water quality variables simultaneously. Unlike single-task models that require training separate models for each target variable [32, 53, 57], the multi-task approach leverages shared learning of hidden relationships among variables, improving predictive performance, and reducing computational costs, which is critical for modeling complex biogeochemical processes [23].
@@
Trustworthiness evaluation framework
@@
Generalization to unseen basins. To test the model’s generalizability, we employed a spatial holdout strategy in which 80% of the basins were randomly selected for training and the remaining 20% were for testing. This ensures that the test basins are completely unseen during training, allowing us to evaluate the model’s ability to generalize to new geographic regions characterized by distinct geographic, climatic, and hydrological conditions.
Data sources
Streamflow and water quality data were extracted from the U.S. Geological Survey (USGS) National Water Information System (NWIS) database (https://waterdata.usgs.gov/nwis). Meteorological variables were extracted from the gridMET dataset (https://www.climatologylab.org/gridmet.html). Vegetation indices were acquired from the Global Land Surface Satellite (GLASS) dataset (http://www.glass.umd.edu/Download.html). Rainfall chemistry data were retrieved from the National Atmospheric Deposition Program/National Trends Network (NADP/NTN) (https://nadp.slh.wisc.edu/networks/national-trends-network/). Basin attributes were obtained from the Geospatial Attributes of Gages for Evaluating Streamflow, version II (GAGES-II) database (https://www.sciencebase.gov/catalog/item/631405bbd34e36012efa304a). Processed data for the 482 basins used in this study are publicly available at https://figshare.com/s/e0151c12b6e6482bae83.
Code and data availability
Python scripts for downloading raw data are available at: https://github.com/fkwai/geolearn/tree/master/hydroDL/data. The LSTM modeling code developed in this study is provided in the supplementary file. Python (3.8.18) codes for statistical analysis and plotting are available from the authors upon request.
@@ Acknowledgements
Fig. S6. Boxplot of Kling-Gupta Efficiency (KGE) values for the testing basins under a random spatial training-testing split, evaluated across 20 predicted water quality variables associated with physical/chemical properties, geochemical weathering processes, and nutrient cycling, respectively. Each boxplot shows the median (central line), interquartile range (IQR, represented by the boxes spanning the first (Q1) to the third quartile (Q3)), and whiskers extending to and . The number labeled on the box indicates the median.